# Hydrodynamics of a disk in a thin film of weakly nematic fluid subject to linear friction

### v1

@@ III.1 Velocity field
1 $6{w}^{2}t$ $-8t{w}^{2}\left(3-5{w}^{2}-60t\right)$ $6t{w}^{2}\big{[}10-40{w}^{2}+35{w}^{4}-160\left(3-7{w}^{2}\right)t+13440t^{2}\big{]}$
2 $10{w}^{4}t$ $-12t{w}^{4}\left(5-7{w}^{2}-140t\right)$ $6t{w}^{4}\big{[}35-112{w}^{2}+84{w}^{4}-448\left(5-9{w}^{2}\right)t+80640t^{2}\big{]}$
3 $14{w}^{6}t$ $-16t{w}^{6}\left(7-9{w}^{2}-252t\right)$ $6t{w}^{6}\big{[}84-240{w}^{2}+165{w}^{4}-960\left(7-11{w}^{2}\right)t+295680t^{2}\big{]}$
4 $18{w}^{8}t$ $-20t{w}^{8}\left(9-11{w}^{2}-396t\right)$ $6t{w}^{8}\big{[}165-440{w}^{2}+286{w}^{4}-1760\left(9-13{w}^{2}\right)t+823680t^{2}\big{]}$
@@
Table 1: The principal part of the series expansion of $\Psi_{mn}$ is examined for cases of $m=n+1$, $m=n+2$, and $m=n+3$, with $\Psi_{mn}$ defined by Eq. (34). General expressions applicable to values of $n$ are provided by Eqs. (III.1).
The first terms of $\mathcal{P}\Psi_{mn}$ for $m=n+1$, $m=n+2$, and $m=n+3$ are obtained as
@@
We now define the small dimensionless number
@@
As previously noted, the evaluation of integrals in Eqs. (III.1) is complex. To facilitate analytical advances, we investigate the scenario of weak fluid anisotropy, that is, $|\beta|\ll 1$. Furthermore, we restrict our analysis to cases of $|A|\ll 1$, see Eq. (11). Both $A$ and $\beta$ are signed quantities, allowing for both positive and negative values in a physical context. Hence, the integrands presented in Eqs. (III.1) can be expressed as a multi-Taylor expansion in the small parameters $A$ and $\beta$. This facilitates a straightforward analytical evaluation of the resulting integrals with respect to $\phi$ in Eqs. (III.1), which otherwise is not obvious. Our considerations are limited to terms up to the first order in the multi-Taylor series expansions.
@@
$\bm{v}=V\left(\bm{v}^{\mathrm{IB}}+\bm{v}^{*}\right)\,.$ (38)
In this context, $\bm{v}^{\mathrm{IB}}$ represents the solution corresponding to an isotropic Brinkman fluid of $A=\beta=0$, while $\bm{v}^{*}$ denotes the leading-order correction arising from anisotropy. The former is obtained as [59]
@@ III.2 Pressure field
In the present framework there are no restrictions on the magnitude of $b$. It can be arbitrarily small or large.
@@
$p=\frac{V\nu_{3}}{a}\left(p^{\mathrm{IB}}+p^{*}\right)\,,$ (42)
@@ IV Resistance coefficients
it can be shown that
$F_{\parallel}$ $=a\nu_{3}\int_{0}^{2\pi}\left.\left(\Pi_{\rho\theta}\cos\theta-\Pi_{\rho\rho}\sin\theta+f_{\parallel}\right)\right|_{\rho=a}\mathrm{d}\theta\,,$ (48a)
$F_{\perp}$ $=a\nu_{3}\int_{0}^{2\pi}\left.\left(\Pi_{\rho\rho}\sin\theta+\Pi_{\rho\theta}\cos\theta+f_{\perp}\right)\right|_{\rho=a}\mathrm{d}\theta\,.$ (48b)
@@
$f_{\parallel}$ $=\epsilon_{+}\left(v_{\rho,\rho}\left(3\cos\theta+\cos(3\theta)\right)-\Pi_{\rho\theta}\left(\sin\theta+\sin(3\theta)\right)+\frac{v_{\rho}+v_{\theta,\theta}}{\rho}\left(\cos\theta-\cos(3\theta)\right)\right),$ (49)
$f_{\perp}$ $=\epsilon_{-}\left(v_{\rho,\rho}\left(3\sin\theta-\sin(3\theta)\right)+\Pi_{\rho\theta}\left(\cos\theta-\cos(3\theta)\right)+\frac{v_{\rho}+v_{\theta,\theta}}{\rho}\left(\sin\theta+\sin(3\theta)\right)\right),$ (50)
@@
$v_{\rho}^{\parallel}$ $=\sum_{n=0}^{\infty}g_{n}^{\parallel}(\rho)\cos\big{(}(2n+1)\theta\big{)}\,,$ (52a)
$v_{\theta}^{\parallel}$ $=\sum_{n=0}^{\infty}h_{n}^{\parallel}(\rho)\sin\big{(}(2n+1)\theta\big{)}\,,$ (52b)
$p^{\parallel}$ $=\sum_{n=0}^{\infty}w_{n}^{\parallel}(\rho)\cos\big{(}(2n+1)\theta\big{)}\,$ (52c)
for motion parallel to the director. Similarly, for motion perpendicular to the director, the sine and cosine are interchanged in these expressions, along with the utilization of the radial functions $g_{n}^{\perp}$, $h_{n}^{\perp}$, and $w_{n}^{\perp}$. It follows from the boundary conditions imposed at the surface of the disk that $g_{0}^{\parallel}=g_{0}^{\perp}=h_{0}^{\perp}=1$ and that $h_{0}^{\parallel}=-1$. In addition, it follows from the incompressibility equation that $\mathrm{d}g_{i}^{\parallel}/\mathrm{d}\rho=\mathrm{d}g_{i}^{\perp}/\mathrm{d}\rho=0$ for $i\in\{0,1\}$ at $\rho=a$. Accordingly, the resistance forces are obtained as
$\frac{F_{\parallel}}{\pi a\nu_{3}}$ $=\left.w_{0}^{\parallel}+\frac{\mathrm{d}}{\mathrm{d}\rho}\left(\left(1+\epsilon_{+}\right)h_{0}^{\parallel}+\epsilon_{+}h_{1}^{\parallel}\right)\right|_{\rho=a},$ (53a)
$\frac{F_{\perp}}{\pi a\nu_{3}}$ $=\left.w_{0}^{\perp}-\frac{\mathrm{d}}{\mathrm{d}\rho}\left(\left(1+\epsilon_{-}\right)h_{0}^{\perp}-\epsilon_{-}h_{1}^{\perp}\right)\right|_{\rho=a}.$ (53b)
We now introduce the hydrodynamic resistance coefficients relating the translational velocities to the corresponding components of the resistance force via $F_{\parallel}=-R_{\parallel}V_{\parallel}$ and $F_{\perp}=-R_{\perp}V_{\perp}$. After some algebraic manipulation, which involves expressing $w_{0}^{\parallel}$ and $w_{0}^{\perp}$ in terms of velocity as dictated by the governing equations of fluid motion, it can be shown that
$\frac{F_{\parallel}}{\pi a\nu_{3}}=\left.\left(a\alpha_{\parallel}\right)^{2}+\frac{A}{4}\frac{\mathrm{d}h_{1}^{\parallel}}{\mathrm{d}\rho}+\left(1+\frac{A}{8}\right)\frac{\mathrm{d}^{2}h_{0}^{\parallel}}{\mathrm{d}\rho^{2}}\right|_{\rho=a}$ (54)
@@
$\frac{F_{\perp}}{\pi a\nu_{3}}=\left.\left(a\alpha_{\perp}\right)^{2}+\frac{A}{4}\frac{\mathrm{d}h_{1}^{\perp}}{\mathrm{d}\rho}-\left(1+\frac{A}{8}\right)\frac{\mathrm{d}^{2}h_{0}^{\perp}}{\mathrm{d}\rho^{2}}\right|_{\rho=a}$ (55)
for motion perpendicular to the director. It is noteworthy to mention that Eqs. (54) and (55) do not impose any further restriction on $A$ and $\beta$.
Particularly, for $\beta=0$ we find $h_{0}^{\perp}=-h_{0}^{\parallel}$ and $h_{1}^{\perp}=h_{1}^{\parallel}$. Thus, we infer that, in this particular scenario, the resistance coefficient remains independent of the orientation of the director to the considered leading order in $A$ and $\beta$. Alternatively, it can be shown that the total resistance force exerted on the translating disk can be expressed as
@@
wherein $\mu^{\mathrm{IB}}=1/R^{\mathrm{IB}}$ is the hydrodynamic mobility in an isotropic Brinkman fluid medium. $k_{\parallel}^{A}=R_{\parallel}^{A}/R^{\mathrm{IB}}$ and $k_{\parallel}^{\beta}=R_{\parallel}^{\beta}/R^{\mathrm{IB}}$ are the leading-order corrections resulting from anisotropy. In the limit of ${c}=\alpha_{\parallel}a\to 0$, we obtain to leading order
@@
with $\gamma$ denoting Euler’s constant. Conversely, in the opposite limit of ${c}=\alpha_{\parallel}a\to\infty$, we obtain
@@
Figure 3: Streamlines and quiver plots illustrating the image flow field $\bm{v}^{*}$ induced by a disk translating at speed $V$in a two-dimensional, uniaxially anisotropic Brinkman fluid. The disk moves along the axis of anisotropy into the positive $x$-direction. Results are presented for three values of the parameter $c=\alpha_{\parallel}a$, namely $c=0.1$ (left column), $c=1$ (middle column), and $c=10$ (right column). Additionally, we consider values of the anisotropy parameters $A$ and $\beta$ as $(A,\beta)=(0.5,0)$ (top row) and $(A,\beta)=(0,0.5)$ (bottom row). Analytical solutions are depicted in the top halves ($y>0$), while FEM results are shown in the bottom halves ($y<0$).
@@
Figure 4: Streamlines and quiver plots illustrating the image flow field $\bm{v}^{*}$ for a disk translating perpendicular to the director in a two-dimensional, uniaxially anisotropic Brinkman fluid, here into the positive $y$-direction. Results are displayed for three values of the parameter $c=\alpha_{\parallel}a$, namely for $c=0.1$ (left column), $c=1$ (middle column), and $c=10$ (right column). Moreover, two different combinations of the anisotropy parameters $A$ and $\beta$ are considered, namely $(A,\beta)=(0.5,0)$ (top row) and $(A,\beta)=(0,0.5)$ (bottom row). We depict analytical solutions in the right-hand halves ($x>0$) and FEM results in the left-hand halves ($x<0$).
@@
Figure 5: Again, streamlines and quiver plots illustrate the image flow field $\bm{v}^{*}$ induced by a translating disk in a two-dimensional, uniaxially anisotropic Brinkman fluid. The motion here is parallel to the director in panels (a) and (c), which include theoretical solutions in the top halves ($y>0$) and FEM results in the bottom halves ($y<0$). In panels (b) and (d) motion is perpendicular to the director, which includes theoretical solutions on the right-hand sides ($x>0$) and FEM results on the left-hand sides ($x<0$). We set $c=\alpha_{\parallel}a=1$. Here, the anisotropy parameters $A$ and $\beta$ are set to $(A,\beta)=(-2,0)$ in (a) and (b), while $(A,\beta)=(0,-4)$ in (c) and (d).
@@
In Fig. 1, we illustrate the variations of the correction factors with respect to $c=\alpha_{\parallel}a$, as derived above theoretically (solid lines) and as determined from finite-element simulations (symbols), which we outline below. Squares and circles represent the results obtained for motion parallel and perpendicular to the director, respectively. The simulations validate the equality $k_{\parallel}^{A}=k_{\perp}^{A}$, demonstrating the independence of the drag coefficient on the direction of motion. Furthermore, it is evident that the correction attributed to anisotropy in friction, quantified by $\beta$, outweighs that of anisotropy in viscosity, quantified by $A$. Additionally, we observe that the impact of the anisotropy onto friction when compared to the isotropic case to leading order is more pronounced for motion perpendicular than parallel to the director.

### v2

@@ III.1 Velocity field
1 $6{w}^{2}t$ $-8{w}^{2}t\left(3-5{w}^{2}-60t\right)$ $10{w}^{2}t\big{[}6-24{w}^{2}+21{w}^{4}-96\left(3-7{w}^{2}\right)t+8064t^{2}\big{]}$
2 $10{w}^{4}t$ $-12{w}^{4}t\left(5-7{w}^{2}-140t\right)$ $14{w}^{4}t\big{[}15-48{w}^{2}+36{w}^{4}-192\left(5-9{w}^{2}\right)t+34560t^{2}\big{]}$
3 $14{w}^{6}t$ $-16{w}^{6}t\left(7-9{w}^{2}-252t\right)$ $18{w}^{6}t\big{[}28-80{w}^{2}+55{w}^{4}-320\left(7-11{w}^{2}\right)t+98560t^{2}\big{]}$
4 $18{w}^{8}t$   $-20{w}^{8}t\left(9-11{w}^{2}-396t\right)$   $22{w}^{8}t\big{[}45-120{w}^{2}+78{w}^{4}-480\left(9-13{w}^{2}\right)t+224640t^{2}\big{]}$
@@
Table 1: The principal part of the series expansion of $\Psi_{mn}$ is examined for cases of $m=n+1$, $m=n+2$, and $m=n+3$, with $\Psi_{mn}$ defined by Eq. (35). General expressions applicable to values of $n$ are provided by Eqs. (37). Here, $t=1/\sigma^{2}$.
Defining $t=1/\sigma^{2}$, the first terms of $\mathcal{P}\Psi_{mn}$ for $m=n+1$, $m=n+2$, and $m=n+3$ are obtained as
@@
We now define the dimensionless number
@@
As previously noted, the evaluation of integrals in Eqs. (30) is complex. To facilitate analytical advances, we investigate the scenario of weak fluid anisotropy, that is, $|\beta|\ll 1$. Furthermore, we restrict our analysis to cases of $|A|\ll 1$, see Eq. (11). Both $A$ and $\beta$ are signed quantities, allowing for both positive and negative values in a physical context. Hence, the integrands presented in Eqs. (30) can be expressed as a multi-Taylor expansion in the small parameters $A$ and $\beta$. This facilitates a straightforward analytical evaluation of the resulting integrals with respect to $\phi$ in Eqs. (30), which otherwise is not obvious. Our considerations are limited to terms up to the first order in the multi-Taylor series expansions.
@@
$\boldsymbol{v}=V\left(\boldsymbol{v}^{\mathrm{IB}}+\boldsymbol{v}^{*}\right)\,.$ (39)
In this context, $\boldsymbol{v}^{\mathrm{IB}}$ represents the solution corresponding to an isotropic Brinkman fluid of $A=\beta=0$, while $\boldsymbol{v}^{*}$ denotes the leading-order correction arising from anisotropy. The former is obtained as [89]
@@ III.2 Pressure field
In the present framework there are no restrictions on the magnitude or sign of $b$. It can be arbitrarily small or large.
@@
$p=\frac{V\nu_{3}}{a^{2}}\left(p^{\mathrm{IB}}+p^{*}\right)\,,$ (43)
@@ IV Resistance coefficients
the frictional forces can be expressed as
$F_{\parallel}$ $=a\nu_{3}\int_{0}^{2\pi}\left.\left(\Pi_{\rho\theta}\cos\theta-\Pi_{\rho\rho}\sin\theta+{\pi_{\parallel}}\right)\right|_{\rho=a}\mathrm{d}\theta\,,$ (49a)
$F_{\perp}$ $=a\nu_{3}\int_{0}^{2\pi}\left.\left(\Pi_{\rho\rho}\sin\theta+\Pi_{\rho\theta}\cos\theta+{\pi_{\perp}}\right)\right|_{\rho=a}\mathrm{d}\theta\,.$ (49b)
@@
${\pi_{\parallel}}$ $=\epsilon_{+}\left(v_{\rho,\rho}\left(3\cos\theta+\cos(3\theta)\right)-\Pi_{\rho\theta}\left(\sin\theta+\sin(3\theta)\right)+\frac{v_{\rho}+v_{\theta,\theta}}{\rho}\left(\cos\theta-\cos(3\theta)\right)\right),$ (50)
${\pi_{\perp}}$ $=\epsilon_{-}\left(v_{\rho,\rho}\left(3\sin\theta-\sin(3\theta)\right)+\Pi_{\rho\theta}\left(\cos\theta-\cos(3\theta)\right)+\frac{v_{\rho}+v_{\theta,\theta}}{\rho}\left(\sin\theta+\sin(3\theta)\right)\right),$ (51)
@@
$v_{\rho}^{\parallel}$ $={V_{\parallel}}\sum_{n=0}^{\infty}g_{n}^{\parallel}(\rho)\cos\big{(}(2n+1)\theta\big{)}\,,$ (53a)
$v_{\theta}^{\parallel}$ $={V_{\parallel}}\sum_{n=0}^{\infty}h_{n}^{\parallel}(\rho)\sin\big{(}(2n+1)\theta\big{)}\,,$ (53b)
$p^{\parallel}$ $={\frac{V_{\parallel}\nu_{3}}{a^{2}}}\,\sum_{n=0}^{\infty}w_{n}^{\parallel}(\rho)\cos\big{(}(2n+1)\theta\big{)}\,$ (53c)
for motion parallel to the director. Similarly, for motion perpendicular to the director, the sine and cosine are interchanged in these expressions, along with the utilization of the radial functions $g_{n}^{\perp}$, $h_{n}^{\perp}$, and $w_{n}^{\perp}$. In particular, for $\beta=0$, we find that $g_{n}^{\perp}=(-1)^{n}g_{n}^{\parallel}$ and $h_{n}^{\perp}=(-1)^{n+1}h_{n}^{\parallel}$; however, $w_{n}^{\perp}=(-1)^{n}w_{n}^{\parallel}$ only when also $b=0$. It follows from the boundary conditions imposed at the surface of the disk that $g_{0}^{\parallel}=g_{0}^{\perp}=h_{0}^{\perp}=1$ and that $h_{0}^{\parallel}=-1$. In addition, it follows from the incompressibility equation that $\mathrm{d}g_{i}^{\parallel}/\mathrm{d}\rho=\mathrm{d}g_{i}^{\perp}/\mathrm{d}\rho=0$ for $i\in\{0,1\}$ at $\rho=a$. Accordingly, the resistance forces are obtained as
$\frac{F_{\parallel}}{\pi\nu_{3}V_{\parallel}}$ $=\left.w_{0}^{\parallel}+a\,\frac{\mathrm{d}}{\mathrm{d}\rho}\left(\left(1+\epsilon_{+}\right)h_{0}^{\parallel}+\epsilon_{+}h_{1}^{\parallel}\right)\right|_{\rho=a},$ (54a)
$\frac{F_{\perp}}{\pi\nu_{3}V_{\perp}}$ $=\left.w_{0}^{\perp}-a\,\frac{\mathrm{d}}{\mathrm{d}\rho}\left(\left(1+\epsilon_{-}\right)h_{0}^{\perp}-\epsilon_{-}h_{1}^{\perp}\right)\right|_{\rho=a}.$ (54b)
We now introduce the hydrodynamic resistance coefficients relating the translational velocities to the corresponding components of the resistance force via $F_{\parallel}=-R_{\parallel}V_{\parallel}$ and $F_{\perp}=-R_{\perp}V_{\perp}$. After some algebraic manipulation, which involves expressing $w_{0}^{\parallel}$ and $w_{0}^{\perp}$ in terms of velocity as dictated by the governing equations of fluid motion, we obtain
$\frac{-R_{\parallel}}{\pi\nu_{3}}=\left.\left(a\alpha_{\parallel}\right)^{2}+a\,\frac{A}{4}\frac{\mathrm{d}h_{1}^{\parallel}}{\mathrm{d}\rho}+a^{2}\left(1+\frac{A}{8}\right)\frac{\mathrm{d}^{2}h_{0}^{\parallel}}{\mathrm{d}\rho^{2}}\right|_{\rho=a}$ (55)
@@
$\frac{-R_{\perp}}{\pi\nu_{3}}=\left.\left(a\alpha_{\perp}\right)^{2}+a\,\frac{A}{4}\frac{\mathrm{d}h_{1}^{\perp}}{\mathrm{d}\rho}-a^{2}\left(1+\frac{A}{8}\right)\frac{\mathrm{d}^{2}h_{0}^{\perp}}{\mathrm{d}\rho^{2}}\right|_{\rho=a}$ (56)
for motion perpendicular to the director. We mention that Eqs. (55) and (56) in their given form do not impose any further restriction on the magnitude of the values of $A$ and $\beta$. We assume small magnitudes of $A$ and $\beta$ to derive approximate expressions for $h_{0}^{\perp}$, $h_{0}^{\parallel}$, $h_{1}^{\perp}$, and $h_{1}^{\parallel}$. These expressions, in turn, provide leading-order approximations for the resistance coefficients.
Particularly, for $\beta=0$, we have $h_{0}^{\perp}=-h_{0}^{\parallel}$ and $h_{1}^{\perp}=h_{1}^{\parallel}$. Thus, we infer that, in this particular scenario, the resistance coefficient remains independent of the orientation of the director to the considered leading order in $A$ and $\beta$. Alternatively, the total resistance force exerted on the translating disk can be expressed as
@@
wherein $\mu^{\mathrm{IB}}=1/R^{\mathrm{IB}}$ is the hydrodynamic mobility in an isotropic Brinkman fluid medium. $k_{\parallel}^{A}=R_{\parallel}^{A}/R^{\mathrm{IB}}$ and $k_{\parallel}^{\beta}=R_{\parallel}^{\beta}/R^{\mathrm{IB}}$ are the leading-order corrections resulting from anisotropy. In the limit of ${c}=\alpha_{\parallel}a\to 0$, hence, $\alpha_{\perp}=(1-\beta)\alpha_{\parallel}\to 0$ due to Eq. (38), we obtain to leading order
@@
with $\gamma$ denoting Euler’s constant. Physically, this limit corresponds to a situation of weak friction with the substrate or the surrounding environment. Conversely, in the opposite limit of ${c}=\alpha_{\parallel}a\to\infty$, hence, $\alpha_{\perp}=(1-\beta)\alpha_{\parallel}\to\infty$, corresponding to a scenario of significant friction with the substrate or the surrounding environment, we find
@@
Figure 4: Streamlines and quiver plots illustrating the image flow field $\boldsymbol{v}^{*}$ induced by a disk translating at speed $V$ in a two-dimensional, uniaxially anisotropic Brinkman fluid. The disk moves along the axis of anisotropy into the positive $x$-direction. Results are presented for three values of the parameter $c=\alpha_{\parallel}a$, namely $c=0.1$ (left column), $c=1$ (middle column), and $c=10$ (right column). Additionally, we consider values of the anisotropy parameters $A$ and $\beta$ as $(A,\beta)=(0.5,0)$ (top row) and $(A,\beta)=(0,0.5)$ (bottom row). Analytical solutions are depicted in the top halves ($y>0$), while FEM results are shown in the bottom halves ($y<0$).
@@
Figure 5: Streamlines and quiver plots illustrating the image flow field $\boldsymbol{v}^{*}$ for a disk translating perpendicular to the director in a two-dimensional, uniaxially anisotropic Brinkman fluid, here into the positive $y$-direction. Results are displayed for three values of the parameter $c=\alpha_{\parallel}a$, namely for $c=0.1$ (left column), $c=1$ (middle column), and $c=10$ (right column). Moreover, two different combinations of the anisotropy parameters $A$ and $\beta$ are considered, namely $(A,\beta)=(0.5,0)$ (top row) and $(A,\beta)=(0,0.5)$ (bottom row). We depict analytical solutions in the right-hand halves ($x>0$) and FEM results in the left-hand halves ($x<0$).
@@
Figure 6: Streamlines and quiver plots illustrate the image flow field $\boldsymbol{v}^{*}$ induced by a translating disk in a two-dimensional, uniaxially anisotropic Brinkman fluid. The motion here is parallel to the director in panels (a) and (c), which include theoretical solutions in the top halves ($y>0$) and FEM results in the bottom halves ($y<0$). In panels (b) and (d) motion is perpendicular to the director, which includes theoretical solutions on the right-hand sides ($x>0$) and FEM results on the left-hand sides ($x<0$). We set $c=\alpha_{\parallel}a=1$. Here, the anisotropy parameters $A$ and $\beta$ are set to $(A,\beta)=(-2,0)$ in (a) and (b), while $(A,\beta)=(0,-4)$ in (c) and (d).
@@
In Fig. 2, we illustrate the variations of the correction factors with respect to $c=\alpha_{\parallel}a$, as derived above theoretically (solid lines) and as determined from finite-element simulations (symbols), which we outline below. Squares and circles represent the results obtained for motion parallel and perpendicular to the director, respectively. The simulations validate the equality $k_{\parallel}^{A}=k_{\perp}^{A}$, demonstrating the independence of the drag coefficient on the direction of motion. Furthermore, it is evident that the correction attributed to anisotropy in friction, quantified by $\beta$, outweighs that of anisotropy in viscosity, quantified by $A$. This behavior is far from trivial. One conceivable reason lies with the nature of two-dimensional fluid flows themselves. Generally, in an isotropic fluid, the Green’s function quantifying the flows induced by a locally applied net force diverges logarithmically with the distance from the point of force application [91, 92]. It is the linear friction with the environment that regularizes this divergence in Brinkman fluids so that the induced flows decay towards zero with infinite distance [22]. This subtle role stresses the importance of the effect of friction in general for two-dimensional fluid flows. Additionally, we observe that the impact of the anisotropy onto friction when compared to the isotropic case to leading order is more pronounced for motion perpendicular than parallel to the director, at least in an intermediate regime. As inferred from Eq. (62), the sign of the hydrodynamic mobility is determined by the expression $1-k_{i}^{A}A-k_{i}^{\beta}\beta$, where $i\in\{\parallel,\perp\}$. Given that $A$ and $\beta$ are assumed to be much less than 1, this quantity is always positive.